OpenStudy (anonymous):

If the sin 30° = , then the cos 60° = _____.

OpenStudy (anonymous):

you didnt say anything,you just put = sign

OpenStudy (anonymous):

\[\sin 30=t\] \[\cos 60=\sin (90-60)=\sin 30=t\]

OpenStudy (anonymous):

because SIN A = COS(90 - A) COS A = SIN(90 - A)

OpenStudy (anonymous):

Oh sorry! it's = 1/2

OpenStudy (anonymous):

@jdoe0001